OpenStudy (anonymous):

giving medals for answers and help! Find the midpoint between the following pairs of points: (5,7),(-3,5)

OpenStudy (anonymous):

can I just plot this?@jdoe0001

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 5}}\quad ,&{\color{blue}{7}})\quad &({\color{red}{ -3}}\quad ,&{\color{blue}{ 5}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2}\right)\)

OpenStudy (anonymous):

0,5?

OpenStudy (anonymous):

@ @jdoe0001

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 5}}\quad ,&{\color{blue}{7}})\quad &({\color{red}{ -3}}\quad ,&{\color{blue}{ 5}}) \end{array}\qquad \left(\cfrac{{\color{red}{ -3}} + {\color{red}{ 5}}}{2}\quad ,\quad \cfrac{{\color{blue}{ 5}} + {\color{blue}{ 7}}}{2}\right)\)

OpenStudy (anonymous):

o hold on ill figre

OpenStudy (anonymous):

1,6? @jdoe0001

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thanks :) now ill know how to do that :)

OpenStudy (jdoe0001):

yw